OpenStudy (anonymous):

x^2/9 - (y-1)^2/16 = 1 name the vertices, and foci

OpenStudy (muhammad_nauman_umair):

let me revise it

OpenStudy (anonymous):

ok thank you :)

OpenStudy (muhammad_nauman_umair):

a=9 b=16 As centre is (0,1) c^2=a^2+b^2 then c=+5and -5 as hyperbola is along x-axis so, vertices are (0+5,1) and(0-5,1) i.e. (5,1),(-5,1)

OpenStudy (anonymous):

is (0,1) the focii?

OpenStudy (muhammad_nauman_umair):

no it is centre

OpenStudy (anonymous):

What is the foci?

OpenStudy (muhammad_nauman_umair):

(5,1) and (-5,1) are focci i mistakenly wrote it as vertices

OpenStudy (muhammad_nauman_umair):

now i vertices are

OpenStudy (muhammad_nauman_umair):

(9,1),(-9,1)

OpenStudy (sirm3d):

@Muhammad_Nauman_Umair \(a^2=9\), not \(a=9\)

OpenStudy (muhammad_nauman_umair):

oh :( i need to go to sleep i m again mistaken @sirm3d u r write a should b 3 and b should be 4

OpenStudy (muhammad_nauman_umair):

right not write again mistaken @sirm3d